OpenStudy (anonymous):

a boat moves 6 kilometers upstream in the smae amount of time it moves 20 kilometers downstream. if the rate of the current is 5 kilometers per hour, find the rate of the boat in still water. a. 9 2/7 kilometers per hour b. 2 1/7 kilometers per hour c. 9 2/7 kilometers per hour d. 8 4/7 kilometers per hour

OpenStudy (amistre64):

s = d/t s = 6/t and s = 20/t

OpenStudy (amistre64):

b + r = 20/t b - r = 6/t might be a better scenario....

OpenStudy (amistre64):

b + 5 = 20 b - 5 = 6 b = 25 and b = 11 25x = 11x ?? cant be right... :) close, but not right

OpenStudy (anonymous):

d = rt 6 = (x - 5)t for the upstream 20 = (x + 5)t for downstream Solve both for t 6/(x - 5) and 20/(x + 5) Since the time is the same, put these equal to each other 6/(x - 5) = 20(x + 5) cross multiply 6(x + 5) = 20(x - 5) distribute 6x + 30 = 20x - 100 Solve 130 = 14x x = 130/14 = 65/7 = 9 2/7 kph